# Microsoft Research Montréal, Canada

### v1

@@ 3.3 Evaluation during model development
LLM benchmarks are widely used to compare the quality and safety progress made by new model releases, resulting in model leaderboards on different scenarios. The Stanford HELM [191] leaderboard for example shows the performance of different LLM models on benchmarks, and these benchmarks include societal impact and bias-related measures. Their HELM (‘holistic framework for evaluating foundation models’) framework [129] uses scenarios, and measures seven metrics. Those are accuracy, calibration, robustness, efficiency, but also more social impact-oriented fairness, bias, and toxicity. Each scenario focuses on one use cases, and consists of a dataset of instances, such as the LegalBench set of legal reasoning tasks [89], or medical board exam problem sets [107]. The larger BIG-bench (“Beyond the Imitation Game benchmark”) [21] consists of 200+ tasks, contributed by hundreds of authors at a variety of institutes. More specific benchmarks for trustworthiness such as DecodingTrust, in turn focus on subsets such as toxicity, stereotyping, adversarial and out-of distribution robustness, privacy, machine ethics, and fairness  [222],
Paradoxically, while these benchmarks include aspects of societal impacts, bias and as toxicity, these benchmarks do not necessarily cover the aspects that matter most in a specific application context. Benchmarks are generally geared towards structured comparisons between models, not towards evaluating end-user applications in practice. This means that they may not be particularly suitable for a specific application and the people involved in its usage. In addition, using such large benchmarks can be quite resource intensive, making ‘lite’ versions necessary that are less comprehensive. Both Helm and BIG-Bench are also implemented as Lite versions. However, the evaluation differences that arise from specific, lighter implementations of benchmarks can significantly impact model comparison results [189]. This makes it necessary to go beyond these benchmarks, and ensure suitable evaluations for the application at hand to avoid deriving conclusions about safety or responsibility devoid from actual application concerns.
@@ 3.3.2 Combining IR and generative AI evaluation metrics
It is challenging that societal impact measure standards, including for bias, fairness and other types of impact are yet scarce for IR product settings. For example, Smith et al. [186] provide an overview of different metrics available for recommendation bias and fairness evaluation and the challenges practitioners face when choosing between them. In some cases, it may be more appropriate to for example focus on ‘traditional’ performance and accuracy metrics, but understand the performance and subsequent quality of experiences for different groups of people by segmenting/slicing results between groups. This then assumes the possibility of being able to define groups, or relies on more advanced methods to find clusters that may - or may not- be actionably different in performance.
Specific methods might also be necessary to match new techniques. For example, Retrieval Augmented Generation (RAG) might be used to include more reliable information in a specific domain, and reduce hallucination in a LLM setting. However, RAG does not necessarily fully solve every hallucination related issue. Specific frameworks that fit an application context are still necessary to evaluate these techniques and their actual impact on aspects such as factuality within that context. One example is Saad-Falcon et al. [173], who present an evaluation framework, ARES, for RAG-assisted Question & Answering settings. This framework uses three evaluation scores: context relevance of the retrieved information, answer faithfulness - the answer’s grounding in the retrieved context, and answer relevance to the question asked. These are similar to IR-evaluations, but might need adjustment to the setting at hand, and used datasets need to reflect actual needs in current circumstances.
@@ 3.3.3 LLMs to evaluate LLM
Beyond specific metrics, ongoing research is investigating the efficacy of LLMs to evaluate LLMs [224, 181]. For example, [224] et al. use an LLM to rate the factuality of a long-form response to prompts, while additionally also using Google Search. While promising, such more complex evaluation constellations also lead to additional complexity in understanding what is being evaluated, and changes therein as the evaluator LLM changes. This leads to having to validate the validation in itself [181]. While a human-and-LLM agent collaboration can help in this validation (as in e.g. [181]’s EvalGen approach), the evaluation criteria cannot be fully separated from observation of model outputs, resulting in a feedback loop from output to adjusted evaluation criteria.
@@ 3.4 Evaluation pre/post system release
Whether evaluations are done online or offline can deeply impact results. Offline evaluations - even when using thoughtful standards - might not reflect what actual end-users do in real-life settings, or system performance over time. Online evaluations similarly are limited to what is in scope in for example included metrics, but include actual interactions. It involves field testing; getting an IR system online and out to actual users and analyzing their interactions with the system. It can include methods such as controlled experiments or extended A/B testing, and analysis of interactions; Hoffmann provides an overview of most common techniques used in IR settings [104].
@@ 3.4.2 Stress testing, red-teaming and qualitative end-user evaluations
Other types of stress testing, and assessment of security issues, as well as evaluation of experiences of actual users is necessary. A helpful discussion is provided by Khlaaf, who points out the need for careful consideration what methods and terminology are appropriate for evaluations that probe for vulnerabilities of a system towards the outside world, including for example red teaming for safety issues that can arise from using a system [116].
Beyond metrics and quantitative analysis oriented methods, it is crucial to apply a combination of safety/security-inspired methods, user design and UX research methods to understand the actual reactions of users. For example, red team approaches in which “a group of people authorized and organized to emulate a potential adversary’s attack or exploitation capabilities against an enterprise’s security posture” (following the definition from the National Institute of Standards and Technology , NIST) in themselves also lead to design questions. For example, using LLMs as red teamers themselves  [162] by generating risky prompts at scale, or using large-scale human red teaming efforts with thousands of participants who need access points might yield different results. Similarly, model characteristics might have consequences on red teaming results. Ganguli et al. [73], for instance find that RLHF models are increasingly difficult to red team as they scale, while they don’t find similar challenges for other models. Interestingly this means that techniques such RLHF (reinforcement learning from human feedback) which are explicitly meant to help align agents with human preferences could also result in challenges in evaluating the systems that use them.
@@ 3.5 Societal impact of a system beyond its direct implementation and use
The impact of system usage can have impact much beyond the direct usage context. For example, the increasing demand for data and compute power of LLMs has environmental impact. Such impact can be hard to calculate, even while methods have been developed in both the IR and LLM communities around reducing environmental harm[176] and sustainability industry teams exist to ensure more energy efficient data centers for both environmental as well as monetary reasons. Others in turn try to assess whether LLMs could help in generating more green code, and develop metrics to assess the code’s ‘green capacity’ based on earlier sustainability metrics [208].
@@ 3.6 Sharing evaluation methods
From the above selection of methods, which is no mean is comprehensive, it is clear that practitioners have to carefully pick and choose which methods work for them. However, different organizations come from different evaluation traditions.
Incentives to share methods and results might not align with practical product team incentives and pressures. Metrics and standards for evaluations from actual practice are often not shared in scientific literature. Security community-styled (external) red and (internal) blue teams, Trust & Safety incident monitoring approaches, IR-communities’ existing offline and online user feedback methods, or UX product testing approaches might be more (or less) top of mind depending on the organization and prior expertise. This means there is a gap in our generative IR literature in terms of shared understanding of actual practices and efficacy of methods  [52]. If we as a community are to properly address the social risks as outlined in 2.2.1, it is imperative we find fast and effective ways to share these methods and align them with practical needs. Especially with the increasing speed of the field, the variety of fields involved, and volume of new techniques.
@@ 4.2.1 Organizational Factors
For the above arrangements, getting to collections of concrete examples of what has worked in the past is increasingly important. AI developments are speeding up, and increasingly diverse professional communities are both being impacted and getting involved. This makes efficient and effective coordination even more important. For policy makers, governmental agencies and journalists it may be hard to get an overview of which professional communities can provide actionable advice - especially with new AI developments being ‘louder’ than, for example, long-standing IR communities. Inside of companies, in order to benefit from external advice or research, tech teams still have to navigate how to best work with external organizations. Researchers and non-governmental organizations in turn have to know where to invest their time and expertise most effectively, and how to offer actionable advice to appropriate individuals or teams in tech companies. This includes big picture scenario planning of where to best invest, and how to create incentives that truly will have a positive impact. Implicit hierarchies of the value of different types of produced knowledge (e.g. ’being the first’ or ’more technically complex’), but also a simple lack of knowledge about how certain processes work, can stand in the way of sharing of paved paths towards desired results, and of sharing these in accessible ways. It can also involve very pragmatic on-the-ground work, such as knowing how to set up contractual arrangements that work for all parties (not a skill commonly taught in IR or AI-related programs).
@@ 4.2.2 Data-focused methods
For effective action, however, it is crucial that generative IR researchers who care about mitigating negative societal impact have access to ways to learn how to effectively organize and navigate organizational and political structures. Implicit hierarchies in what knowledge is appreciated in generative IR circles can become a hurdle in effectively identifying and addressing the risks outlined in earlier sections, §2.2.1, §2.2.2, and §2.2.3. A critical factor is knowing which concrete situations matter, what to ask for in those situations and how to assess whether impacts and risks are successfully steered.
@@ 5 Conclusion
In this chapter we have presented a discussion on the sociotechnical implications of generative AI for information access. Our deliberations are grounded in how these emerging technologies are currently being applied in IR applications as well as their future applications as being envisioned by practitioners and researchers. It is important to recognize that our sociotechnical visions of what information access should look like in the future is not just shaped by what emerging technologies like generative AI make plausible, but our collective visions for the future of information access should in turn shape AI technologies themselves. Mitra [143] proposed the hierarchy of IR stakeholder needs shown in 2 and argued that IR research and system development require a fundamental shift towards re-centering societal needs and that we should reimagine information access as a vehicle for radically new emancipatory futures for our society. When contemplating the implications of emerging technologies, we risk of falling in the trap of limiting ourselves to how the technology (and its process of development) is today, rather than how it can be or should be in the future. Neither generative AI nor its application in the context of information access is predetermined. So, while it is important that we consider potential harms of contemporary applications of generative AI in the context of information access, we close with some open question for the reader: If not this status quo, then what and how? What is the future of information access that we want to imagine for our collective wellbeing and liberation, and how can generative AI be another tool in the toolbox towards that transformation?

### v2

@@ 3.3 Evaluation during model development
LLM benchmarks are widely used to compare the quality and safety progress made by new model releases, resulting in model leaderboards on different scenarios. The Stanford HELM [200] leaderboard for example shows the performance of different LLM models on benchmarks, and these benchmarks include societal impact and bias-related measures. Their HELM (‘holistic framework for evaluating foundation models’) framework [132] uses scenarios, and measures seven metrics. Those are accuracy, calibration, robustness, efficiency, but also more social impact-oriented fairness, bias, and toxicity. Each scenario focuses on one use case, and consists of a dataset of instances, such as the LegalBench set of legal reasoning tasks [91], or medical board exam problem sets [110]. The larger BIG-bench (“Beyond the Imitation Game benchmark”) [22] consists of 200+ tasks, contributed by hundreds of authors at a variety of institutes. More specific benchmarks for trustworthiness such as DecodingTrust, in turn focus on subsets such as toxicity, stereotyping, adversarial and out-of distribution robustness, privacy, machine ethics, and fairness [231], while for example the much more specific recurring TREC Fair Ranking track competitively evaluates systems according to how fairly they rank documents on a specific test task [65].
Paradoxically, while these benchmarks include aspects of societal impacts such as bias and toxicity, they do not necessarily cover the aspects that matter most in a specific application context in practice. Benchmarks are generally geared towards structured comparisons between models, not towards evaluating end-user applications in practice. This means that they may not be particularly suitable for a specific application and the people involved in its usage. In addition, using such large benchmarks can be quite resource intensive, making ‘lite’ versions necessary that are less comprehensive. Both Helm and BIG-Bench are also implemented as Lite versions. However, the evaluation differences that arise from specific, lighter implementations of benchmarks can significantly impact model comparison results [198]. This makes it necessary to go beyond these benchmarks, and ensure suitable evaluations for the application at hand to avoid deriving conclusions about safety or responsibility devoid from actual application concerns.
@@ 3.3.2 Combining IR and generative AI evaluation metrics
It is challenging that standards for measuring societal impact, including bias, fairness and etc. are yet scarce in IR product settings. For example, Smith et al. [195] provide an overview of different metrics available for evaluating bias and fairness in recommendation systems, and the challenges practitioners face when choosing between them. In some cases, it may be more appropriate to for example focus on ‘traditional’ performance and accuracy metrics, but study the performance and subsequent quality of experiences for different groups of people by segmenting/slicing results by group. This approach assumes the ability to define relevant groups, or relies on more advanced methods to find clusters that may—or may not—have significant differences in performance or quality.
Specific methods might also be necessary to match new techniques. For example, Retrieval Augmented Generation (RAG) might be used to include more reliable information in a specific domain and reduce hallucinations in an LLM setting. However, RAG does not necessarily fully solve every hallucination-related issue. Specific frameworks that fit an application context are still necessary to evaluate these techniques and their actual impact on aspects such as factuality within that context. One example is Saad-Falcon et al. [181], who present an evaluation framework, ARES, for RAG-assisted Question & Answering settings. This framework uses three evaluation scores: context relevance of the retrieved information, answer faithfulness (the answer’s grounding in the retrieved context), and answer relevance to the question asked. These are similar to IR-evaluations, but might need adjustment to the setting at hand, and datasets used need to reflect actual needs in current circumstances.
@@ 3.3.3 LLMs to evaluate LLM
Beyond specific metrics, ongoing research is investigating the efficacy of LLMs to evaluate LLMs (LLM-as-judge)[233, 190]. For example, [233] et al. use an LLM to rate the factuality of a long-form response to prompts, while also using Google Search. While promising, such more complex evaluation constellations also lead to additional complexity in understanding what is being evaluated, and changes therein as the evaluator LLM changes. This leads to having to validate the validation in itself [190]. While a human-and-LLM agent collaboration can help in this validation (as in e.g. [190]’s EvalGen approach), the evaluation criteria cannot be fully separated from observation of model outputs, resulting in a feedback loop from output to adjusted evaluation criteria.
@@ 3.4 Evaluation pre/post system release
Whether evaluations are done online or offline can deeply impact results. Offline evaluations—even when using thoughtful standards—might not reflect what actual end-users do in real-life settings, or system performance over time. Online evaluations similarly are limited to which metrics have been instrumented and how actual user interactions are captured. It involves field testing; getting an IR system online and out to actual users and analyzing their interactions with the system. It can include methods such as controlled experiments or extended A/B testing, and analysis of interactions; Hoffmann provides an overview of most common techniques used in IR settings [106].
@@ 3.4.2 Stress testing, red-teaming and qualitative end-user evaluations
Beyond metrics and quantitative analysis oriented methods, it is crucial to apply a combination of safety/security-inspired methods, user design and UX research methods to understand the actual reactions of users.
The logistics around red teaming can provide a good glimpse into the importance of appropriate combinations of methods. Red teaming is a common way to test LLM applications for undesirable system responses [140, 149]. Red teaming can be automated using for example sets of (generated) prompts, or done in full by human red teamers, including both the general public, or invited experts. Using LLMs as red teamers [169] by generating risky prompts at scale, or using large-scale human red teaming efforts with thousands of participants who need access points, might yield different results. Human red team approaches in which “a group of people authorized and organized to emulate a potential adversary’s attack or exploitation capabilities against an enterprise’s security posture” (if we follow the definition from the National Institute of Standards and Technology, NIST) also lead to questions about tooling, recruiting and operational process design. Markov et al. [140] for example provide a helpful discussion of practical data challenges in content moderation use cases. In turn, model characteristics might have consequences on red teaming results. Ganguli et al. [75], for instance find that RLHF (reinforcement learning from human feedback) models are increasingly difficult to red team as they scale, while they don’t find similar challenges for other models. Interestingly this means that techniques such RLHF that are explicitly meant to help align agents with human preferences could also result in challenges in evaluating the systems that use them.
This means that like any evaluation method, red teaming has to be combined with other types of stress testing, assessment of security issues, as well as evaluation of experiences of actual users. Khlaaf points out the need for carefully considering what methods and terminology are appropriate for evaluations that probe for vulnerabilities of a specific system towards the outside world [119].
@@ 3.5 Societal impact of a system beyond its direct implementation and use
The impact of a system can reach much beyond its direct usage context. For example, the increasing demand for data and compute power of LLMs has environmental impact. However, such indirect impact can be hard to calculate without deep expertise. It is crucial to spend the time to evaluate evaluations methods for their suitability. Methods have been developed in both the IR and LLM communities around reducing environmental harm[184] and sustainability industry teams exist to ensure more energy efficient data centers for both environmental as well as monetary reasons. Others in turn try to assess whether LLMs could help in generating more green code, and develop metrics to assess the code’s ‘green capacity’ based on earlier sustainability metrics [217].
@@ 3.6 Sharing evaluation methods
From the above selection of methods, which is by no means comprehensive, it is clear that practitioners have to carefully pick and choose which methods work for them. However, different organizations come from different evaluation traditions.
Incentives to share methods and results might not align with practical product team incentives and pressures. Metrics and standards for evaluations from actual practice are often not shared in scientific literature. Security community-style (external) red and (internal) blue teams, Trust & Safety incident monitoring approaches, IR-communities’ existing offline and online user feedback methods, or UX product testing approaches might be more (or less) top of mind depending on the organization and prior expertise. This means there is a gap in the generative IR literature in terms of shared understanding of actual practices and efficacy of methods [53]. If we as a community are to properly address the social risks as outlined in 2.2.1, it is imperative we find fast and effective ways to share these methods and align them with practical needs. Especially with the increasing speed of the field, the variety of fields involved, and volume of new techniques.
@@ 4.2.1 Organizational Factors
For the above arrangements, getting to collections of concrete examples of what has worked in the past is increasingly important. AI developments are speeding up, and increasingly diverse professional communities are both being impacted and getting involved. This makes efficient and effective coordination even more important. For policy makers, governmental agencies and journalists it may be hard to get an overview of which professional communities can provide actionable advice—especially with new AI developments being ‘louder’ than, for example, long-standing IR communities. Inside of companies, in order to benefit from external advice or research, tech teams still have to navigate how to best work with external organizations. Researchers and non-governmental organizations in turn have to know where to invest their time and expertise most effectively, and how to offer actionable advice to appropriate individuals or teams in tech companies. This includes big picture scenario planning of where to best invest, and how to create incentives that truly will have a positive impact. Implicit hierarchies of the value of different types of produced knowledge (e.g. ’being the first’ or ’more technically complex’), but also a simple lack of knowledge about how certain processes work, can stand in the way of sharing of paved paths towards desired results, and of sharing these in accessible ways. It can also involve very pragmatic on-the-ground work, such as knowing how to set up contractual arrangements that work for all parties (not a skill commonly taught in IR or AI-related programs).
@@ 4.2.2 Data-focused methods
For effective research-informed mitigations, however, it is crucial that generative IR researchers have access to ways to learn how to effectively organize and navigate organizational and political structures, or how to communicate their results to others. Implicit hierarchies in what knowledge is appreciated in generative IR circles can become a hurdle in effectively identifying and addressing the risks outlined in earlier sections, §2.2.1, §2.2.2, and §2.2.3. A critical factor is knowing which concrete situations matter, what to ask for in those situations and how to assess whether impacts and risks are successfully steered.
@@ 5 Conclusion
In this chapter we have presented a discussion on the sociotechnical implications of generative AI for information access. These deliberations are grounded in how these emerging technologies are currently being applied in IR applications as well as their future applications as being envisioned by practitioners and researchers. It is important to recognize that sociotechnical visions of what information access should look like in the future are not just shaped by what emerging technologies like generative AI make plausible, but that visions for the future of information access in turn shape AI technologies themselves. Mitra [150] proposed the hierarchy of IR stakeholder needs shown in 2 and argued that IR research and system development require a fundamental shift towards re-centering societal needs and that we should reimagine information access as a vehicle for alternative futures. When contemplating the implications of emerging technologies, we risk of falling in the trap of limiting ourselves to how the technology (and its process of development) is today, rather than how it can be or should be in the future. Neither generative AI nor its application in the context of information access is predetermined. So, while it is important that we consider potential harms of contemporary applications of generative AI in the context of information access, we close with some open question for the reader: If not this status quo, then what—and especially how? What is the future of information access that we want to imagine for our collective wellbeing, and how can generative AI be another tool in the toolbox towards that transformation?